# Population Transformer: Learning Population-level Representations of Neural Activity

### v2

@@ Appendix A Architectures and training
Pretrained PopT The core Population Transformer consists of a transformer encoder stack with 6 layers, 8 heads. All layers ($N=6$) in the encoder stack are set with the following parameters: $d_{h}=512$, $H=8$, and $p_{\text{dropout}}=0.1$. We pretrain the PopT model with the LAMB optimizer (You et al., 2019) ($lr=1e-4$), with a batch size of $n_{\text{batch}}=256$, and train/val/test split of 0.98, 0.01, 0.01 of the data. We pretrain for 500,000 steps, and record the validation performance every 1,000 steps. Downstream evaluation takes place on the weights with the best validation performance. We use the intermediate representation at the [CLS] token $d_{h}=512$ and put a linear layer that outputs to $d_{out}=1$ for fine-tuning on downstream tasks. These parameters for pretraining were the same for any PopT that needed to be pretrained (hold-one-out subject, subject subsets, ablation studies).
@@
Downstream Training For both PopT models, we train with these parameters: AdamW optimizer (Loshchilov & Hutter, 2017), $lr=5e^{-4}$ where transformer weights are scaled down by a factor of 10 ($lr_{t}=5e^{-5}$), $n_{batch}=256$, a Ramp Up scheduler (ildoonet, 2024) with warmup 0.025 and Step LR gamma 0.99, reducing 100 times within the 2000 total steps that we train for. For Linear and DeepNN models, we train with these parameters: AdamW optimizer (Loshchilov & Hutter, 2017), $lr=5e^{-4}$, $n_{batch}=256$, a Ramp Up scheduler (ildoonet, 2024) with warmup 0.025 and Step LR gamma 0.95, reducing 25 times within the 17,000 total steps we train for. For all downstream decoding, we use a fixed train/val/test split of 0.8, 0.1, 0.1 of the data.
@@ Appendix B Model and Compute Requirements
Table 4: Parameter counts. Since PopT takes existing temporal embeddings as input, the number of parameters that must be trained is an order of magnitude less than recent end-to-end approaches.
GPU count GPU type Time to train TFLOPS
@@
Table 5: Pretraining compute requirements Based on published train times (none were given for LaBraM) it is evident that PopT has smaller hardware and shorter training time requirements.
@@ Appendix D Data
1 19 91 Thor: Ragnarok 1.83
Fantastic Mr. Fox 1.75
The Martian 0.5 x
2 12 100 Venom 2.42
@@
Guardians of the Galaxy 2.5
Guardians of the Galaxy 2 3
Avengers: Infinity War 4.33
Black Panther 1.75
Aquaman 3.42 x
3 18 91 Cars 2 1.92 x
Lord of the Rings 1 2.67
Lord of the Rings 2 (extended edition) 3.92
4 9 135 Megamind 2.58
Toy Story 1.33
Coraline 1.83 x
5 11 205 Cars 2 1.75 x
@@
6 12 152 Incredibles 1.15
Shrek 3 1.68 x
Megamind 2.43
7 6 109 Fantastic Mr. Fox 1.5
8 4.5 72 Sesame Street Episode 1.28
9 16 102 Ant Man 2.28
10 12 173 Cars 2 1.58 x
Spider-Man: Far from Home 2.17
@@
Table 6: Subject statistics Subjects used in PopT training, and held-out downstream evaluation. Table taken from Wang et al. (2022). The second column shows the number of uncorrupted, electrodes that can be Laplacian re-referenced. The average amount of recording data per subject is 4.3 (hrs).
Appendix E Interpretation Methods
@@
Connectivity analysis We start with a pretrained PopT. To test a particular channel’s contribution to connectivity, we mask it with all zeros. Then, we consider the remaining unmasked channels and ask, how much increase do we see in the pretraining channel-wise loss? Recall that this objective is to determine whether or not a channel has had its inputs swapped with random activity. If the increase in loss is large, then we infer that the masked channel provided important context for this task. Using this delta as a measure for connectivity, we then average across the Desikan-Killiany atlas regions (Alexander et al., 2019) and produce a plot using mne-connectivity (Gramfort et al., 2013).
@@
Scaled Attention Weight First, we obtain an attention weight matrix across all trials which includes weights between all tokens. Then, we perform attention rollout (Abnar & Zuidema, 2020) across layers to obtain the contributions of each input channel by the last layer. We take the resulting last layer of rollout weights for all channels, where the target is the [CLS] token, normalize within subject, and scale by ROC AUC to obtain the Scaled Attention Weight per channel. Finally, we plot the 0.75 percentile weight per region, as mapped by the Destrieux atlas (Destrieux et al., 2010) using Nilearn (contributors, ).
Appendix F Connectivity
@@
Figure 10: Full connectivity for our 7 test subjects. We compare between traditional connectivity analaysis performed via coherence (top row in each section) and the analysis based on our PopT pretrained weights (bottom row in each section). We note that our analysis usually recovers the strongest points of connectivity fromt the traditional analysis. Coherence was computed using scikit-learn’s (Pedregosa et al., 2011) signal.coherence.
Appendix G Frozen scaling
@@
Figure 11: Pretraining is critical to frozen PopT performance that scales with the number of channels. As in Figure 3, we see that pretraining results in better downstream decoding and better scaling with the number of added channels. However, unlike in Figure 3, the PopT weights are frozen during fine-tuning, and only the linear classification head is updated. Bands show standard error across subjects. Results are shown for a frozen PopT with BrainBERT inputs.
Appendix H Frozen ablation
Sentence onset Speech/Non-speech Pitch Volume
Frozen PopT $\mathbf{0.73}\pm\mathbf{0.06}$ $\mathbf{0.72}\pm\mathbf{0.08}$ $0.59\pm 0.06$ $\mathbf{0.63}\pm\mathbf{0.07}$
w/o cls $0.67\pm 0.08$ $0.68\pm 0.07$ $0.58\pm 0.04$ $0.60\pm 0.07$
w/o replace loss $0.69\pm 0.07$ $0.69\pm 0.09$ $\mathbf{0.59}\pm\mathbf{0.06}$ $0.62\pm 0.06$
w/o position encoding $0.70\pm 0.07$ $0.69\pm 0.07$ $0.56\pm 0.08$ $0.61\pm 0.06$
Table 7: An ablation study of the components of our approach for the frozen PopT. During pretraining, we alternate using either only the CLS or token contrastive component of the loss. We fine-tune these weights on all subjects. We find that both components contribute to the full model’s performance.

### v4

@@ Appendix A Architectures and training
Pretrained PopT The core Population Transformer consists of a transformer encoder stack with 6 layers, 8 heads. All layers in the encoder stack are set with the following parameters: $d_{h}=512$, $H=8$, and $p_{\text{dropout}}=0.1$. We pretrain the PopT model with the LAMB optimizer (You et al., 2019) ($lr=5e-4$), with a batch size of $n_{\text{batch}}=256$, and train/val/test split of 0.89, 0.01, 0.10 of the data. We pretrain for 500,000 steps, and record the validation performance every 1,000 steps. Downstream evaluation takes place on the weights with the best validation performance. We use the intermediate representation at the [CLS] token $d_{h}=512$ and put a linear layer that outputs to $d_{out}=1$ for fine-tuning on downstream tasks. These parameters for pretraining were the same for any PopT that needed to be pretrained (across temporal embeddings, hold-one-out subject, ablation studies).
[…]
Pretraining task: Channel-wise pretraining The token level objective is to determine whether a channels activity has been swapped with activity from a random time. Precisely, activity from each channel $i$ is drawn from a time $t_{i}$. All channels are drawn from the same time $t_{i}=t$, and then 10% of the channels are randomly selected to have their activity replaced with activity from a randomly selected channel, taken from a random point in time $t_{i}\neq t$. Then, the channel-wise outputs, $\tilde{y}_{i}\in\mathbb{R}^{d}$, of the Population Transformer are passed through a linear layer to obtain scalar predictions $\hat{y}_{i}$. The objective function is the BCE between these predictions and the labels $y^{*}_{i}$: $\mathcal{L}_{C}=\frac{1}{|S_{A}|+|S_{B}|}\sum_{i}y^{*}_{i}\log(p(\hat{y}))+(1-y^{*}_{i})\log(p(\hat{y}_{i}))$ where $y^{*}_{i}=\mathbf{1}(t_{i}\neq t)$. Then, the complete pretraining objective is $\mathcal{L}=\mathcal{L}_{C}+\mathcal{L}_{N}$
@@
Downstream Training For PopT models, we train with these parameters: AdamW optimizer (Loshchilov & Hutter, 2017), $lr=5e^{-4}$ where transformer weights are scaled down by a factor of 10 ($lr_{t}=5e^{-5}$), $n_{batch}=128$, a Ramp Up scheduler (ildoonet, 2024) with warmup 0.025 and Step LR gamma 0.95, reducing 100 times within the 2000 total steps that we train for. For Linear and DeepNN models, we train with these parameters: AdamW optimizer (Loshchilov & Hutter, 2017), $lr=1e^{-3}$, $n_{batch}=256$, a Ramp Up scheduler (ildoonet, 2024) with warmup 0.025 and Step LR gamma 0.95, reducing 100 times within the 17,000 total steps we train for. For all downstream decoding, we use a fixed train/val/test split of 0.8, 0.1, 0.1 of the data. For non-BrainBERT models, we provide configuration files for downstream decoding parameters in our codebase.
@@ Appendix B Model and Compute Requirements
BIOT (Yang et al., 2024) 3M
Table 4: Parameter counts. Since PopT takes existing temporal embeddings as input, the number of trainable parameters is an order of magnitude less than some recent end-to-end approaches.
# GPUs GPU type Time to train TFLOPS
@@
BIOT (Yang et al., 2024) 8 NVIDIA RTX A6000 (48G)
Table 5: Pretraining compute requirements Based on published train times (none were given for LaBraM and BIOT) it is evident that PopT has smaller hardware and shorter training time requirements.
@@ Appendix D Dataset details
1 19 91 Thor: Ragnarok 2.07
Fantastic Mr. Fox 1.91
The Martian 2.90 x
2 12 100 Venom 2.60
@@
Guardians of the Galaxy 2.66
Guardians of the Galaxy 2 3.00
Avengers: Infinity War 3.73
Black Panther 1.85
Aquaman 3.52 x
3 18 91 Cars 2 1.90 x
Lord of the Rings 1 2.94
Lord of the Rings 2 (extended edition) 4.06
4 12 152 Incredibles 1.31
Shrek 3 1.87 x
Megamind 1.75
5 6 109 Fantastic Mr. Fox 1.54
6 9 135 Megamind 0.81
Toy Story 1.32
Coraline 1.6 x
7 11 205 Cars 2 1.67 x
@@
8 4.5 121 Sesame Street Episode 1.41
9 16 72 Ant Man 1.0
10 12 173 Cars 2 1.57 x
Spider-Man: Far from Home 2.33
@@
Table 6: Subject statistics Subjects used in BrainBERT training, and held-out downstream evaluation. The number of uncorrupted, electrodes that can be Laplacian re-referenced are shown in the second column. The average amount of recording data per subject is 5.55 (hrs).
Appendix E Hold out subject pretraining generalizability
Figure 10: Gains in decoding performance are available to new subjects even on TOTEM pretrained PopT. Same experiment as Figure 6 but with TOTEM embedding.
Appendix F Random electrode ensemble performance
Figure 11: Downstream decoding performance on random electrode subsets. To check if our original channel ensemble ordering inflated performance, we perform downstream decoding on 3 randomly generated electrode ensembles. The random electrode ensembles perform roughly similar to our reported values, with the exception of a few low-electrode count ensembles for Sentence Onset. These exceptions may be due to strong decodability of Sentence Onset at specific electrodes. Shaded bands show the standard error across subjects.
Appendix G Interpretation Methods
@@
Connectivity analysis We start with a pretrained PopT. To test a particular channel’s contribution to connectivity, we omit it from the input (more details in Appendix I). Then, we consider the remaining unmasked channels and ask: how does this change the pretraining channel-wise loss? Recall that this objective is to determine if a channel has had its inputs swapped with random activity. If the change in loss is large, we infer that the masked channel provided important context. Using the magnitude of this delta as a measure for connectivity, we then average across the Desikan-Killiany regions (Desikan et al., 2006) and produce a plot using mne-connectivity (Gramfort et al., 2013).
@@
Scaled Attention Weight First, we obtain an attention weight matrix across all trials which includes weights between all tokens. Then, we perform attention rollout (Abnar & Zuidema, 2020) across layers to obtain the contributions of each input channel by the last layer. We take the resulting last layer of rollout weights for all channels, where the target is the [CLS] token, normalize within subject, and scale by ROC AUC to obtain the Scaled Attention Weight per channel. Finally, we take the 0.75 percentile score per DKT region (Klein & Tourville, 2012), average across 5 fine-tuning runs, and plot using Nilearn (contributors, ).
Appendix H Functional brain region comparison
Figure 12: Comparison of functional maps as identified by our method vs traditional measures. Scaled Attention Weight vs Fraction of Significant Electrodes per DKT (Klein & Tourville, 2012) region for the Speech vs. Non-speech task. Fraction of significant electrodes from Wang et al. (2024). The Pearson’s $r$ correlation coefficient is 0.4 between the two metrics. Error bars are standard deviation across 5 fine-tuning runs.
Appendix I Connectivity
Figure 13: Schematic of connectivity analysis To determine the influence of some channel, $i$, on another channel $j$, we first measure the baseline performance of the pretrained PopT on the replace-only objective. Then, we omit $i$ from the input, and measure how the performance on the channel-wise objective is perturbed on $j$. See also Algorithm 1.
Algorithm 1 Connectivity measurement between channels $i$ and $j$
$j<i$, $x\in\mathbb{R}^{N_{C}\times d}$triangleright $N_{C}$ is the number of channels, $d$ is the embedding dimension.
$\hat{y}_{\text{baseline}}\leftarrow P(x)$triangleright $P$ is a pretrained PopT, $\hat{y}_{\text{baseline}}\in\mathbb{R}^{N_{C}}$
while $n\leq N_{\text{samples}}$ do
     $x_{\text{omitted}}\leftarrow\text{Concat}(x[:i],x[i+1:])$triangleright Remove the $i^{th}$ channel from the input
     $\hat{y}_{\text{perturbed}}\leftarrow P(x_{\text{omitted}})$
     $\text{Influence}=|\hat{y}_{\text{baseline}}-\hat{y}_{\text{perturbed}}|$triangleright How much did the prediction change?
     $\text{AvgConnectivity}\leftarrow\text{AvgConnectivity}+\text{Influence}[j]/n$
end while
Figure 14: Electrode level connectivity. Connectivity between all channels for the same subject shown in Figure 8. Outliers at the 2-percentile are snapped to color map floor and ceiling.
@@ Appendix I Connectivity
Figure 15: Region connectivity for test subjects. Continued from Figure 8; this figures shows the rest of the test subjects. We compare between traditional connectivity analysis performed via coherence (top row in each section) and the analysis based on our PopT pretrained weights (bottom row in each section). We note that our analysis usually recovers the strongest points of connectivity from the traditional analysis. Coherence was computed using scikit-learn’s (Pedregosa et al., 2011) signal.coherence.
Figure 16: Electrode connectivity for test subjects. Continued from Figure 14; this figures shows the rest of the test subjects. Order is given as in Figure 15.
Subject Correlation
Subject 1 0.42
Subject 2 0.66
Subject 3 0.54
Subject 4 0.55
Subject 6 0.44
Subject 7 0.44
Subject 10 0.50
Table 7: Pearson’s $r$ correlation coefficients between connectivity matrices for test subjects shown in Table 7 and Figure 16.
Appendix J Additional Ablations
Pitch Volume Sent. Onset Speech/Non-speech
PopT $\mathbf{0.74}\pm\mathbf{0.03}$ $\mathbf{0.87}\pm\mathbf{0.03}$ $\mathbf{0.90}\pm\mathbf{0.01}$ $\mathbf{0.93}\pm\mathbf{0.02}$
PopT w/ gaussian blur $0.73\pm 0.03$ $0.86\pm 0.03$ $0.88\pm 0.02$ $0.91\pm 0.03$
PopT w/o channel randomize $0.71\pm 0.02$ $0.85\pm 0.03$ $0.89\pm 0.01$ $0.91\pm 0.03$
Table 8: PopT additional ablation study. We pretrain additional variations of PopT to see their effect on downstream decoding. In ‘PopT w/ gaussian blur’ we fuzz the input coordinate values with a Gaussian, $\mathcal{N}(\mu=0,\sigma=5)$, before position encoding. We hypothesized that augmenting the coordinates during training would help the model generalize better, but no improvements were shown. ‘PopT w/o channel randomize’ replaces a channel with a channel’s own activity at another time as part of the channel-wise pretraining task. We hypothesized that this would help the model identify the channel’s specific variability across time, but no improvements were shown. Shown are ROC-AUC mean and standard error across subjects evaluated at 90 electrodes.
Appendix K Frozen ensemble scaling
@@
Figure 17: Pretraining is critical to frozen PopT performance that scales with the number of channels. Transformer weights are frozen; only the linear classification head is updated during fine-tuning. As in Figure 3, we see that pretraining results in better downstream decoding and scales with the number of added channels. Bands show standard error across subjects.